### v1

@@ Abstract
We propose a procedure for the numerical approximation of invariance equations arising in the moment matching technique associated with reduced-order modeling of high-dimensional dynamical systems. The Galerkin residual method is employed to find an approximate solution to the invariance equation using a Newton iteration on the coefficients of a monomial basis expansion of the solution. These solutions to the invariance equations can then be used to construct reduced-order models. We assess the ability of the method to solve the invariance PDE system as well as to achieve moment matching and recover a system’s steady-state behaviour for linear and nonlinear signal generators with system dynamics up to $n=1000$ dimensions.
@@
reduced-order modeling, moment matching, invariance equations, global polynomial approximation
@@ 1 Introduction
Complex large-scale dynamical systems are ubiquitous in control engineering. They provide a natural framework to model multi-agent systems, energy networks, and coupled systems of electrical signals, to name a few [25]. The high dimensionality of these models, often in the order of thousands or millions of states, introduces significant computational challenges when trying to analyse the behaviour of such systems, a phenomenon know as the “curse of dimensionality” [9]. The curse of dimensionality naturally appears in control systems design where it is often the case that the resulting control law has a complexity that is proportional, or worse, to the complexity of the system being controlled. This occurs in state feedback and observer design, where the controller is of proportional dimension to the system; model predictive control, where an optimization problem is run on the model at every time step to determine a cost-minimizing control input, and adaptive control, where different parameters need to be estimated while constructing a control input. Such control methods may be infeasible or costly to implement in practice when applied directly to the high-dimensional model. Therefore, it is often beneficial to reduce the complexity of the model by making approximations that make the problem easier and more tractable to solve, while maintaining accuracy and/or particular properties of the system.
@@
Reduced-order modeling aims to alleviate the computational complexity of a problem while preserving important characteristics, creating a smaller model that approximates the behaviour of a complex system to some prescribed degree of accuracy. Reduced-order modeling finds application in the analysis and design of systems in many fields, for example mechanical and electrical systems, and weather forecasting [11, 2]. Broadly speaking, model order reduction methods are categorized as being based upon the singular value decomposition, such as Balanced Truncation  [27, 26, 38] and Proper Orthogonal Decomposition [28, 29, 42], or as being based upon the notion of Krylov projection/interpolation, such as moment matching [4, 3], see [1] for a comprehensive overview. Other approaches for model order reduction include energy-based methods [16, 37], reduction around a limit cycle [17], and transfer function interpolation for polynomial systems [10].
@@
In this article, we focus on nonlinear model reduction by moment matching for interconnected systems involving a signal generator and a nonlinear system, building on the theory of output regulation. For recent surveys of techniques in reduced-order modeling by moment matching see e.g. [33, 36]. For linear systems, the concept of moment matching was shown to be in one-to-one relation with the solution to a Sylvester equation and an associated manifold induced by the interconnection of the system and the signal generator [15]. With additional stability assumptions the invariant manifold is attractive, and the solution of the Sylvester equation can further linked to the steady-state response [21]. This relationship was enhanced for nonlinear systems in [20], where the notion of invariance was associated with the solution of a “Sylvester-like” high-dimensional system of nonlinear coupled PDEs. With additional (local) stability assumptions, the invariant manifold associated to the solution of the PDE system is (locally) attractive and, similarly to the linear case, can be related to the notion of steady-state response. This enables the definition of moment matching to be decoupled from the inherently linear concept of frequency response, and [4] generalized the notion of moment matching to nonlinear systems by considering the invariant manifold induced by interconnection of nonlinear systems which is the solution to a PDE system of the form studied in this article.
The interconnection-based interpretation of moment matching presented in [4] leads to many further developments in the theory of moment matching. For nonlinear systems this includes, for example, the two-sided nonlinear moment matching framework [19], a data-driven approach for estimation of the moment of a nonlinear system by fitting a linear combination of basis functions to observed steady-state trajectories [35, 31], the parametrization of all differential-algebraic systems that achieve moment matching [39, 40], and the closed-loop moment matching framework that allows for the relaxation of the stability requirement on the system to be interpolated in the time domain [30]. A recent survey on interconnection-based model order reduction is provided in [36].
The notion of invariance commonly arises in control theory and the analysis of dynamical systems, so PDE systems of the form considered here are ubiquitous and not unique to the problem of model reduction. The output regulation problem for nonlinear control systems requires the solution of a system of PDEs corresponding to an invariance condition [12]. Control design via immersion and invariance was introduced in [5], where the solution to the PDE system is an immersion of the dynamics of the system-to-be-controlled into those of a “target system” having desired closed-loop behaviour, and the solution to the PDE system is used to construct a controller which achieves this closed-loop behaviour in steady-state.
@@
At this point, the main limitation of the proposed method is that it requires numerical approximation of a large-scale system of coupled nonlinear PDEs over a domain that is determined by the dimension of the signal generator. Overalll, for an $n-$dimensional dynamical system and a $d-$dimensional signal generator, we need to solve a system of $nd$ PDEs in $\mathbb{R}^{d}$. In this paper, we focus on low-dimensional signal generators with $d=2$ and high-dimensional dynamics with $n$ up to 1000, a scale for which the use of traditional grid-based methods for PDEs is unpractical. Since the first-order system of nonlinear PDEs arising in the model reduction is similar to a system of Hamilton-Jacobi-Bellman PDEs from optimal feedback control, we propose a numerical scheme based on Newton’s method in conjunction with a global polynomial ansatz for each PDE solution. This approach first explored in the context of Hamilton-Jacobi-Bellman PDEs in [8, 7] and later extended to high-dimensional settings in [23, 22, 6], mitigates the curse of dimensionality, scaling to large-scale systems at a moderate computational cost. The use of orthogonal polynomials for spectral and pseudospectral weighted residual methods for moment matching has been explored in [14] with linear signal generators and low-dimensional dynamics.
@@ 1.1 Contribution
This paper focuses on the development of computational methods for nonlinear moment matching. The main contributions are listed here:
  • We develop and implement a Galerkin method to find approximate solutions to the invariance PDE system associated with the steady-state response of complex systems, and use these solutions to construct a reduced-order model to achieve moment matching.
  • While many theoretical descriptions of reduced-order modeling by moment matching exist, there is a gap in the literature in developing computational methods for approximating the invariance equation for nonlinear moment matching. This paper represents the first attempt at explicitly developing a computational method to achieve these goals. In particular, the method’s ability to handle dynamics with thousands of states is an important novel contribution.
  • The method’s efficacy is demonstrated in different complex examples, including systems of dimension up to 1000, systems where the nonlinearity is sinusoidal rather than polynomial in nature, and interconnected systems involving implicitly defined nonlinear signal generators.
  • The method works in many different problems whenever the invariance PDE arises. In particular, this shows that the method is not only applicable to model order reduction problems, but also to other general control problems, for example output regulation, and immersion and invariance.
@@ 1.2 Organization
The paper is organized as follows. In Section 2 we give an overview of the problem setup and we show how moment matching is defined and how the invariance PDE system arises. Section 3 introduces a Galerkin type method for the numerical computation of an approximate solution $\pi(\cdot)$ to the invariance PDE system. Section 4 reports different numerical experiments assessing the accuracy of the proposed method and its effectiveness in generating a reduced-order model.
@@ 2 Moment Matching for Nonlinear Dynamical Systems
We consider dynamical systems with a state $x(t)\in\mathbb{R}^{n}$ and an output $y(t)\in\mathbb{R}^{p}$ of the form
@@
where the input to the system is given by $u(t)\in\mathbb{R}^{m}$. We assume that the mappings $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{n}$ and $h:\mathbb{R}^{n}\rightarrow\mathbb{R}^{p}$ are $C^{r}$ with $r\geq 2$ and defined locally around the origin. Without loss of generality, we also assume that (1) has an equilibrium at the origin, that is, $f(0,0)=0$ and $h(0)=0$.
@@
where $v(t)\in\mathbb{R}^{m}$ is the output of the signal generator system, and the mappings $s:\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $\ell:\mathbb{R}^{d}\to\mathbb{R}^{m}$ are $C^{r}$ with $r\geq 2$. We also assume that (2) has an equilibrium at the origin.
@@
The system (1) is said to be observable if for different initial conditions $x_{a}(0)\neq x_{b}(0)$, the output trajectories satisfy $h(x_{a}(t))\neq h(x_{b}(t))$ for some $t\in\mathbb{R}$, i.e. the output trajectories are not identical.
@@
Let $\mathcal{R}(x(0),T)$ be the set (reachable set) containing all points $\overline{x}$ for which there exists an input $u$ such that the evolution of (1) from $x(0)\in\mathcal{X}$ satisfies $x(t)\in\mathcal{X}$ for $0\leq t\leq T$ with $x(T)=\overline{x}$. The system (1) is said to be (locally) accessible if for all $x(0)\in\mathcal{X}$ the set $\bigcup_{t\leq T}\mathcal{R}(x(0),t)$ contains a non-empty open subset of $\mathcal{X}$ for all $T>0$.
@@
A point $\omega_{0}$ is said to be Poisson stable if, for each time $T>0$ and each neighbourhood $U^{0}$ about $\omega_{0}$, the flow $\omega^{*}(t)$ of $s(\omega)$ with the initial condition $\omega_{0}$ passes through $U^{0}$ at points $t_{1}>T$ and $t_{2}<-T$.
@@
Clearly, if the signal generator is neutrally stable on a domain $W^{0}$, then $\omega(t)$ is persistent in time and does not decay to zero as time tends towards infinity, for any initial condition $\omega_{0}\in W^{0}$. Furthermore, this implies that the linear approximation of the system around the origin $S=\left.\left[\frac{\partial s}{\partial\omega}\right]\right|_{\begin{subarray}{c}\omega=0\end{subarray}}$ has purely imaginary eigenvalues, and that these eigenvalues are simple.
@@ 2.1 Notion of Moment
(3) $\dot{\omega}(t)$ $=s(\omega(t)),$
$\dot{x}(t)$ $=f(x(t),\ell(\omega(t))),$
$y(t)$ $=h(x(t)).$
The notion of the time-domain moment for a nonlinear dynamical system comes from analysing the steady-state output of the interconnected system given by (3) and makes use of the centre manifold theory. In order to help in our analysis of this interconnected dynamical system, we use the following definitions from [20] and [13].
Definition 2.5 (Centre Manifold, [13]).
A mapping $\pi$ between two coordinates $\omega$ and $x$ is an invariant manifold for a system of differential equations if the solution of the system given by $(\omega(t),x(t))$ with initial condition $(\omega_{0},\pi(\omega_{0}))$, lies on the curve $(\omega(t),\pi(\omega(t)))$.
Furthermore, if $\pi(0)=0$ and the tangent space $\frac{\partial\pi}{\partial\omega}(0)$ is equal to the center subspace of the interconnected system, then it is also called a centre manifold. This subspace corresponds to exactly those eigenvalues of the interconnected system that have zero real part.
Under certain assumptions for the signal generator (2) and the system (1), we can assume that a locally attractive centre manifold exists between the two systems. We therefore make the following assumptions about the functions, based on [20]:
@@
    The system (1) is asymptotically stable in the first approximation about the equilibrium at the origin. That is, the matrix $A=\left.\left[\frac{\partial f}{\partial x}\right]\right|_{\begin{subarray}{c}x=0\\ u=0\end{subarray}}$ has only eigenvalues with negative real part.
Based on these assumptions, we can conclude the following based on [20, Proposition 8.1.1].
@@
If Assumptions (A1) and (A2) hold, then we are guaranteed the existence of a mapping $\pi:\mathbb{R}^{d}\to\mathbb{R}^{n}$ that is defined on a neighbourhood $W^{0}$ about the origin and characterizes a locally attractive centre manifold for (3). This mapping $\pi$ satisfies, for all $\omega\in W^{0}$,
(4) $\frac{\partial\pi}{\partial\omega}s(\omega)$ $=f(\pi(\omega),\ell(\omega)),$
@@
This centre manifold is represented as =$\mathcal{M}=\{(x,\omega):x=\pi(\omega)\}$ for $\pi$ solving (4). For each flow $\omega^{*}(t)$ with $\omega_{0}\in W^{0}$ as an initial condition, the mapping $\pi$ solving (4) produces a well-defined steady-state in the interconnected system $y_{ss}(t)=h(\pi(\omega^{*}(t)))$.
@@
If the initial condition of the signal generator satisfies $\omega_{0}\in W^{0}$ and if the initial condition $x_{0}$ of the system is sufficiently close to the centre manifold, then the centre manifold is said to be locally attractive if we have
[…]
This implies that
(7) $\lim_{t\to\infty}\|y(t)-h(\pi(\omega^{*}(t)))\|=0.$
@@
Assumptions (A1) and (A2) guarantee that (4) has a solution, and that the notion of moment is in one-to-one relation with the steady-state response of output of the interconnected system (3) for initial conditions starting near the centre manifold $\mathcal{M}$. In the rest of the paper, we shall assume that assumptions (A1) and (A2) hold.
Note that in the linear case, if $f$ has a locally asymptotically stable equilibrium, then it is also globally asymptotically stable. Thus, the transient response due to any initial condition $x_{0}$ will disappear, and we can say that the invariant manifold is globally attractive.

### v2

@@ Abstract
We propose a procedure for the numerical approximation of invariance equations arising in the moment matching technique associated with reduced-order modeling of high-dimensional dynamical systems. The Galerkin residual method is employed to find an approximate solution to the invariance equation using a Newton iteration on the coefficients of a monomial basis expansion of the solution. These solutions to the invariance equations can then be used to construct reduced-order models. We assess the ability of the method to solve the invariance PDE system as well as to achieve moment matching and recover the steady-state behaviour of nonlinear systems with state dimension of order 1000 driven by linear and nonlinear signal generators.
@@
reduced-order modeling, moment matching, invariant manifolds, Galerkin methods, global polynomial approximation
@@ 1 Introduction
Complex large-scale dynamical systems are ubiquitous in control engineering, providing a natural framework for modeling multi-agent systems, energy networks, and coupled electrical systems [25]. The high dimensionality of these models, often on the order of thousands or millions of states, introduces significant computational challenges, a phenomenon known as the “curse of dimensionality” [9]. Reduced-order modeling aims to alleviate this complexity while preserving important system characteristics. It finds application across many fields, including mechanical and electrical systems and weather forecasting [11, 2]. Broadly speaking, model order reduction methods are categorized as being based upon the singular value decomposition, such as Balanced Truncation [27, 26, 38] and Proper Orthogonal Decomposition [28, 29, 42], or upon Krylov projection/interpolation, such as moment matching [4, 3]; see [1] for a comprehensive overview. Other approaches include energy-based methods [16, 37], reduction around a limit cycle [17], and transfer function interpolation for polynomial systems [10].
In this paper, we consider potentially high-dimensional systems with a state $x(t)\in\mathbb{R}^{n}$, input $u(t)\in\mathbb{R}^{m}$, and output $y(t)\in\mathbb{R}^{p}$, governed by
$\dot{x}=f(x,u),\qquad y=h(x).$
In particular, we are interested in the case where the input $u$ is generated by an exogenous system with state $\omega(t)\in\mathbb{R}^{d}$, described by
$\dot{\omega}=s(\omega),\qquad v=\ell(\omega).$
Under certain assumptions, the interconnected system obtained by setting $u=v$ admits a mapping $x=\pi(\omega)$ satisfying a system of first-order nonlinear partial differential equations (PDEs) known as the invariance equation, which characterizes an invariant manifold embedding the interconnected system’s dynamics. The goal of this work is to achieve nonlinear model reduction by moment matching by solving this invariance equation, yielding a reduced system with state $r(t)\in\mathbb{R}^{d}$ governed by
$\dot{r}=\bar{f}(r,u),\qquad y_{r}=\bar{h}(r),$
whose output $y_{r}$ approximates the steady-state behaviour of the full-order system under the same input. The precise mathematical setting is given in Section 2.
@@
For recent surveys on moment matching see e.g. [33, 36]. For linear systems, moment matching is in one-to-one relation with the solution to a Sylvester equation induced by the interconnection of the system and the signal generator [15]; with additional stability assumptions the invariant manifold is attractive and links to the steady-state response [21]. This relationship was extended to nonlinear systems in [20], associating invariance with a “Sylvester-like” system of nonlinear coupled PDEs whose (locally) attractive solutions generalize the notion of steady-state response. Building on this, [4] generalized moment matching to nonlinear systems via the invariant manifold induced by the interconnection. Further developments include two-sided nonlinear moment matching [19], data-driven estimation via basis function fitting [35, 31], parametrization of all differential-algebraic interpolants [39, 40], and a closed-loop framework relaxing stability requirements [30]. The notion of invariance also arises beyond model reduction, in output regulation [12] and in immersion and invariance control design [5].
@@
The main limitation of the interconnection-based approach is the need to numerically solve a large-scale system of coupled nonlinear PDEs: for an $n$-dimensional system and a $d$-dimensional signal generator, this amounts to solving $n$ PDEs over $\mathbb{R}^{d}$. In this paper, we propose a numerical scheme based on Newton’s method in conjunction with a global polynomial ansatz for each PDE solution. This approach, first explored for Hamilton-Jacobi-Bellman PDEs in [8, 7] and later extended to high-dimensional settings in [23, 22, 6], mitigates the curse of dimensionality at moderate computational cost. We focus on low-dimensional signal generators with $d=2$ and high-dimensional dynamics with $n$ up to 1000, a scale for which traditional grid-based PDE methods are impractical. The use of orthogonal polynomials for weighted residual methods in moment matching was explored in [14] for linear signal generators and low-dimensional dynamics.
@@ 1.1 Contribution
This paper focuses on the development of computational methods for nonlinear moment matching. The main contributions are as follows:
  • We develop and implement a Galerkin method to find approximate solutions to the invariance PDE system and use these to construct reduced-order models achieving moment matching.
  • The framework is tailored to nonlinear systems driven by low-dimensional signal generators, exploiting the interconnection structure to construct very low-dimensional reduced-order models that accurately reproduce steady-state behaviour and outperform standard model-order reduction (MOR) techniques at the same reduced order.
  • This paper represents the first computational method explicitly developed for approximating the invariance equation in nonlinear moment matching. The method’s ability to handle systems with thousands of states is a key novel contribution.
  • The method’s efficacy is demonstrated on systems of dimension up to 1000, systems driven by non-polynomial signal generator nonlinearities, and interconnected systems with implicitly defined nonlinear signal generators.
  • The method applies beyond model order reduction, to output regulation and immersion and invariance problems.
@@ 1.2 Organization
The paper is organized as follows. In Section 2 we give an overview of the problem setup and we show how moment matching is defined and how the invariance PDE system arises. Section 3 introduces a Galerkin type method for the numerical computation of an approximate solution $\pi(\cdot)$ to the invariance PDE system. Section 4 reports different numerical experiments assessing the accuracy of the proposed method and its effectiveness in generating a reduced-order model. Section 5 discusses the results further and puts them in a wider context.
@@ 2 Moment Matching for Nonlinear Dynamical Systems
We consider dynamical systems with a state $x(t)\in\mathbb{R}^{n}$, an input $u(t)\in\mathbb{R}^{m}$, and an output $y(t)\in\mathbb{R}^{p}$, that can be represented in the form
@@
We assume that the mappings $f:\mathcal{U}\times\mathcal{V}\to\mathbb{R}^{n}$ and $h:\mathcal{U}\to\mathbb{R}^{p}$ are $C^{2}$, defined on open neighbourhoods $\mathcal{U}\subset\mathbb{R}^{n}$ and $\mathcal{V}\subset\mathbb{R}^{m}$ of the origin. We assume that system (1) has an equilibrium and, without loss of generality, that this equilibrium is satisfied at the origin, that is, $f(0,0)=0$ and $h(0)=0$.
@@
where $v(t)\in\mathbb{R}^{m}$ is the output of the signal generator system, and the mappings $s:\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $\ell:\mathbb{R}^{d}\to\mathbb{R}^{m}$ are $C^{2}$. We also assume that (2) has an equilibrium at the origin.
We denote by $\Xi_{t}^{s}:\mathbb{R}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ the flow generated by the vector field $s$, that is, the smooth function with the property that
(3) $\frac{\partial}{\partial t}\Xi_{t}^{s}(\omega_{0})=s(\Xi_{t}^{s}(\omega_{0})),\quad\Xi_{0}^{s}(\omega_{0})=\omega_{0}.$
@@
The system (1) is said to be locally observable at the point $x_{0}$ if for every open neighbourhood $U$ of $x_{0}$, initial conditions $x_{a}(0)=x_{0}\neq x_{b}(0)\in U$, and input $u(\cdot)\in L^{\infty}\left([0,\infty);\mathbb{R}^{m}\right)$ for which the corresponding trajectories $x_{a}(\cdot),x_{b}(\cdot)$ remain in $U$, the output trajectories satisfy $h(x_{a}(t))\neq h(x_{b}(t))$ for some $t>0$, i.e. the output trajectories are not identical.
If we relax the condition and say that this holds when $U$ is the entire space $\mathcal{U}$, then the system (1) is said to be observable at $x_{0}$. If a system is (locally) observable at every $x_{0}$, then it is said to be (locally) observable.
@@
Let $\mathcal{U}\subset\mathbb{R}^{n}$ be a set which contains some open ball with center $x=0$. Let $\mathcal{R}(x(0),T)$ be the set (reachable set) containing all points $\overline{x}$ for which there exists an input $u(\cdot)\in L^{\infty}\left([0,\infty);\mathbb{R}^{m}\right)$ such that the evolution of (1) from $x(0)\in\mathcal{U}$ satisfies $x(t)\in\mathcal{U}$ for $0\leq t\leq T$ with $x(T)=\overline{x}$. The system (1) is said to be (locally) accessible if for all $x(0)\in\mathcal{U}$ the set $\bigcup_{t\leq T}\mathcal{R}(x(0),t)$ contains a non-empty open subset of $\mathcal{U}$ for all $T>0$.
@@
A point $\omega_{0}\in\mathbb{R}^{d}$ is said to be Poisson stable for system (2) if the flow $\Xi_{t}^{s}(\omega_{0})$ generated by the vector field $s(\omega)$ is defined for all $t\in\mathbb{R}$ and, for each neighbourhood $U^{0}$ about $\omega_{0}$ and each $T>0$, there exists $t_{1},t_{2}\in\mathbb{R}$ with $t_{1}>T$ and $t_{2}<-T$, such that the flow satisfies $\Xi_{t_{1}}^{s}(\omega_{0})\in U^{0}$ and $\Xi_{t_{2}}^{s}(\omega_{0})\in U^{0}$.
@@
Clearly, if the signal generator is neutrally stable on a domain $W^{0}$, then $\omega(t)$ is persistent in time and does not decay to zero as time tends towards infinity, for any initial condition $\omega_{0}\in W^{0}$. Furthermore, this implies that the linear approximation of the system around the origin $S=\left.\left[\frac{\partial s}{\partial\omega}\right]\right|_{\begin{subarray}{c}\omega=0\end{subarray}}$ has purely imaginary eigenvalues, and that these eigenvalues are semi-simple.
@@ 2.1 Notion of Moment
(4) $\dot{\omega}$ $=s(\omega),$
$\dot{x}$ $=f(x,\ell(\omega)),$
$y$ $=h(x),$
[…]
The notion of the time-domain moment for a nonlinear dynamical system comes from analyzing the steady-state output of the interconnected system given by (4) and makes use of the centre manifold theory. In order to help in our analysis of this interconnected dynamical system, we first define the notion of an invariant manifold and a centre manifold, which are introduced in more detail in [20], [13], [25].
Definition 2.5 (Invariant Manifold).
Consider a system
(5) $\dot{\xi}=\eta(\xi),$
for $\xi\in\mathbb{R}^{n}$. A set $\mathcal{M}$ is said to be an invariant manifold for (5) if $\xi(0)\in\mathcal{M}\Rightarrow\xi(t)\in\mathcal{M}$ for all $t\in\mathbb{R}$.
Definition 2.6 (Centre Manifold).
Now consider a system of the form
(6) $\dot{\omega}$ $=S\omega+\tilde{s}(\omega,x)$
$\dot{x}$ $=Ax+\tilde{f}(\omega,x),$
where $\omega(t)\in\mathbb{R}^{d},x(t)\in\mathbb{R}^{n}$, $S$ is a $(d\times d)$ matrix having all eigenvalues with zero real part, $A$ is an $(n\times n)$ matrix having all eigenvalues with negative real part, and the functions $\tilde{s}$ and $\tilde{f}$ are $C^{2}$ with the functions and their first order derivatives vanishing at the origin. If $\mathcal{M}=\{(\omega,x):x=\pi(\omega)\}$, for some function $\pi:\mathbb{R}^{d}\to\mathbb{R}^{n}$, is an invariant manifold for (6), it is also called a centre manifold if it is smooth with $\pi(0)=0$ and $\frac{\partial\pi}{\partial\omega}(0)=0$.
Moreover, the centre manifold $\mathcal{M}$ characterized by the mapping $x=\pi(\omega)$ is said to be a locally attractive centre manifold for (6) if there exists a neighbourhood $U^{0}\subset\mathbb{R}^{d}\times\mathbb{R}^{n}$ of $(0,0)$ and constants $C\geq 1,\ \alpha>0$ such that if $(\omega(0),x(0))\in U^{0}$, then
(7) $\|x(t)-\pi(\omega(t))\|\leq Ce^{-\alpha t}\|x(0)-\pi(\omega(0))\|,$
for all $t\geq 0$.
Under certain assumptions for the signal generator (2) and the system (1), we can assume that a locally attractive centre manifold exists between the two systems. We therefore make the following assumptions about the functions:
@@
    The system (1) is locally exponentially stable in the first approximation about the equilibrium at the origin. That is, the matrix $A=\left.\left[\frac{\partial f}{\partial x}\right]\right|_{\begin{subarray}{c}x=0\\ u=0\end{subarray}}$ has only eigenvalues with negative real part.
The following result is a consequence of [20, Proposition 8.1.1]:
@@
Consider the interconnected system (4). If Assumptions (A1) and (A2) hold, then we are guaranteed the existence of a mapping $\pi:W^{0}\subset\mathbb{R}^{d}\to\mathbb{R}^{n}$, defined on a neighbourhood $W^{0}$ about the origin, that characterizes a locally attractive centre manifold for (4). This mapping $\pi$ satisfies, for all $\omega\in W^{0}$,
(8) $\frac{\partial\pi}{\partial\omega}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\omega)}s(\omega)$ $=f(\pi(\omega),\ell(\omega)),$
@@
For each flow $\Xi_{t}^{s}(\omega_{0})$ with initial condition $\omega_{0}\in W^{0}$, the mapping $\pi$ solving (8) produces a well-defined steady-state response in the interconnected system (4), given by
(9) $y_{ss}(t)=h(\pi(\Xi_{t}^{s}(\omega_{0}))).$
Remark 2.8.
@@
Because this centre manifold is locally attractive, (7) implies that for all initial conditions $\omega_{0}\in W^{0}$, if the initial condition $x_{0}$ of the system is sufficiently close to the centre manifold, the output $y(t;\omega_{0},x_{0})$ generated by initial conditions $(\omega_{0},x_{0})$ satisfies
(11) $\lim_{t\to\infty}\|y(t;\omega_{0},x_{0})-h(\pi(\Xi_{t}^{s}(\omega_{0})))\|=0.$
@@
Assumptions (A1) and (A2) guarantee that (8) has a solution, and that the notion of moment is in one-to-one relation with the steady-state output of the interconnected system (4) for initial conditions starting near the centre manifold $\mathcal{M}$. In the rest of the paper, we shall assume that assumptions (A1) and (A2) hold.